Find the product
(x-5)[x-(5+i)][x-(5-i)]

Answer by stanbon(75887)

Find the product
(x-5)[x-(5+i)][x-(5-i)]
----
= (x-5)[(x-5)-i][(x-5)+i]
----
= (x-5)[(x-5)^2+1]
-----
= (x-5)(x^2-10x+26)
-----
= x^3-10x^2-5x^2+26x+50x-130
---
= x^3-15x^2+76x-130
========================
Cheers,
Stan H.
==============